a 100% concentrate is to be mixed with a mixture having a concentrate of 30% to obtain a 50 gallon mixture of 80%. How much of the 100%concentrate will be needed
Answer by mananth(16946)

Price ---------------- quantity
100 % concentrate 100.00% ---------------- x gallons
30% concentrate% 30.00% ------ 50 - x gallons
Mixture 80.00% ---------------- 50
Total 50 gallons
100.00% x + 30.00% ( 50 - x ) = 80.00% * 50
100 x + 30 ( 50 - x ) = 4000
100 x + 1500 - 30 x = 4000
100 x - 30 x = 4000 - -1500
70 x = 2500
/ 70
x = 35.71 gallons 100.00% 100 % concentrate
14.29 gallons 30.00% 30% concentrate%

m.ananth@hotmail.ca